OpenStudy (anonymous):

Find the f '(x). f(x) = 2x^2 ; a=1 Help please!

OpenStudy (cwrw238):

if f(x) = ax^n then f'(x) = anx^(n-1)

OpenStudy (anonymous):

We are not allowed to use the short cut lol.

OpenStudy (anonymous):

@Edwiener What do you mean shortcut?

OpenStudy (anonymous):

The derivative has its long method and short method. We have to use f(x) -f(a)/ x-a

OpenStudy (anonymous):

If so, then mention the formula: Anyway, plug the value a = 1 in, then subtract:

OpenStudy (anonymous):

Sorry, should have mentioned it.

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 f(a) = 2a^2 ------------------------------------------------------- f(x) -f(a) 2x^2 -2a^2 2(x^2 - a^2) 2(x^2 - 1) Do you see where to go from here?

OpenStudy (anonymous):

What happened to the a on the fourth step.

OpenStudy (anonymous):

a = 1, plug it in!

jimthompson5910 (jim_thompson5910):

a = 1, so I replaced 'a' with 1

OpenStudy (anonymous):

Okay, thanks!